let x be a random variable giving the number of heads minus the number of tails in three tosses of a coin. List the elements of the sample space S for the three tosses of the coin and assign a value to each sample point
1
Expert's answer
2019-11-12T09:53:51-0500

The sample space S for the three tosses of the coin is:


"S=\{{HHH,HHT,HTH,HTT,THH,THT,TTH,TTT} \}"

Let "W" be a random variable giving the number of heads minus the number of tails in three tosses of a coin, we assign a value of "\omega" of "W" to each sample point in the following way: 


"\begin{array}{cc}
 \text{Sample points} & & \omega \\
HHH & & 3 \\HHT & & 1 \\HTH & & 1 \\HTT & & -1 \\THH & & 1 \\THT & & -1 \\TTH & & -1\\
 TTT & & -3
\end{array}"

The sample space for "W" is


"S_W=\{{-3, -1, 1, 3} \}"

corresponding to "3T,1H2T,2H1T," and "3H" respectively.


"P(0\ heads\ \&\ 3\ tails)=\binom{3}{0}({1 \over 2})^0({1 \over 2})^3={1 \over 8}"

"P(1\ head\ \&\ 2\ tails)=\binom{3}{1}({1 \over 2})^1({1 \over 2})^2={3 \over 8}"

"P(2\ heads\ \&\ 1\ tail)=\binom{3}{2}({1 \over 2})^2({1 \over 2})^1={3 \over 8}"

"P(3\ heads\ \&\ 0\ tails)=\binom{3}{3}({1 \over 2})^3({1 \over 2})^0={1 \over 8}"

"P(W=-3)={1 \over 8}"

"P(W=-1)={3 \over 8}"

"P(W=1)={3\over 8}"


"P(W=3)={1 \over 8}"